### v2

@@ Abstract
Schrödinger bridge is a diffusion process that steers a given distribution to another in a prescribed time while minimizing the effort to do so. It can be seen as the stochastic dynamical version of the optimal mass transport, and has growing applications in generative diffusion models and stochastic optimal control. In this work, we propose a regularized variant of the Schrödinger bridge with a quadratic state cost-to-go that incentivizes the optimal sample paths to stay close to a nominal level. Unlike the conventional Schrödinger bridge, the regularization induces a state-dependent rate of killing and creation of probability mass, and its solution requires determining the Markov kernel of a reaction-diffusion partial differential equation. We derive this Markov kernel in closed form. Our solution recovers the heat kernel in the vanishing regularization (i.e., diffusion without reaction) limit, thereby recovering the solution of the conventional Schrödinger bridge. Our results enable the use of dynamic Sinkhorn recursion for computing the Schrödinger bridge with a quadratic state cost-to-go, which would otherwise be challenging to use in this setting. We deduce properties of the new kernel and explain its connections with certain exactly solvable models in quantum mechanics.
@@
Schrödinger bridge, Markov kernel, reaction-diffusion PDE, stochastic optimal control.
@@ 1 Introduction
This solution structure was discovered in [56, 57] and later found to hold in more general settings too, e.g., when prior physics is of the form [^1] Here $\bm{f},\bm{g}$ are regular enough so that strong solutions exist for ${\rm{d}}\bm{x}_{t}=\bm{f}(t,\bm{x}_{t}){\rm{d}}t+\sqrt{2}\bm{g}(t,\bm{x}_{t}){\rm{d}}\bm{w}_{t}$.: ${\rm{d}}\bm{x}_{t}=\bm{f}(t,\bm{x}_{t}){\rm{d}}t+\sqrt{2}\bm{g}(t,\bm{x}_{t}){\rm{d}}\bm{w}_{t}$ generalizing Scrödinger’s prior model $\bm{f}=\bm{0},\bm{g}=$ identity. In such generality, one needs to replace the heat kernel $\kappa_{0}$ with the corresponding $(\bm{f},\bm{g})$-dependent Markov kernels associated with the forward and backward Kolmogorov PDEs, which are advection-diffusion PDEs.
@@
Classical SB has been generalized in multiple directions, e.g., with more general drifts and diffusions [42, 12, 47, 43, 61, 49, 48, 64] and with additional bounded domain constraints [11, 26]. However, works on SB with state cost-to-go are limited [65, 23], [46, Ch. 2.6, 2.7]. In particular, investigating SB with quadratic state cost-to-go and deducing closed form kernels for the same, as done here, are novel.
@@ 2 Notations and preliminaries
We will use the following result, which is sometimes referred to as the “central identity” of quantum field theory (QFT) [67, Appendix A].
@@
[69, p. 2][67, p. 15] For a suitably smooth function $f:\mathbb{R}^{n}\mapsto\mathbb{R}$, and matrix $\bm{A}\in{\rm{Sym}}_{++}(n)$,
(9) $\displaystyle\int_{\mathbb{R}^{n}}\!\!\exp\!\left(\!-\frac{1}{2}\bm{x}^{\top}\bm{A}\bm{x}\!\right)\!f(\bm{x}){\rm{d}}\bm{x}=\sqrt{\frac{(2\pi)^{n}}{\det(\bm{A})}}\exp\left(\frac{1}{2}\nabla_{\bm{x}}^{\top}\bm{A}^{-1}\nabla_{\bm{x}}\right)f(\bm{x})\bigg{|}_{\bm{x}=\bm{0}}.$
In particular, for $f(\bm{x})=\exp\langle\bm{b},\bm{x}\rangle$, $\bm{b}\in\mathbb{R}^{n}$, we have $\int_{\mathbb{R}^{n}}\!\exp\!\left(\!-\frac{1}{2}\bm{x}^{\top}\bm{A}\bm{x}+\langle\bm{b},\bm{x}\rangle\!\right)\!{\rm{d}}\bm{x}=\sqrt{\frac{(2\pi)^{n}}{\det(\bm{A})}}\exp\left(\frac{1}{2}\bm{b}^{\top}\bm{A}^{-1}\bm{b}\right)$.
@@ 3.1 Schrödinger factors and reaction-diffusion PDEs
The unique solution to (11) can be determined in terms of the harmonic and coharmonic factors $\varphi,\widehat{\varphi}$, a.k.a. Schrödinger factors, as follows (proof in Appendix A.2). Variants of this result have appeared before; see e.g., [14], [18, Sec. 5].
@@
(15) $-\iota\displaystyle\frac{\partial\widehat{\Psi}}{\partial\tau}=\left(\Delta-q\right)\widehat{\Psi},\qquad\iota\displaystyle\frac{\partial\Psi}{\partial\tau}=\left(-\Delta+q\right)\Psi,\qquad\iota:=\sqrt{-1}.$
The PDEs (15) can be obtained from (12) under imaginary time change (Wick’s rotation) $t\mapsto\iota\tau$, where $q(\cdot)$ plays the role of quantum mechanical potential. Then, $\Psi,\widehat{\Psi}$ are the wave function and its adjoint, and their product $\widehat{\Psi}(t,\cdot)\Psi(t,\cdot)$ returns the probability amplitude as in (14a). In other words, the Schrödinger factors $(\widehat{\varphi},\varphi)$ in Theorem 3.2 are classical analogues of the wave functions, and the relation (14a) is the classical analogue of the celebrated Born’s relation [7]. For further discussions relating (12) and (15), and their solutions, see [2, Ch. 1.4, 2.1].
@@ 3.3 Exactly solvable SBs and closed form kernels
We say that an SB problem is exactly solvable if the corresponding (uncontrolled) Markov kernel $\kappa$ is known in closed form [^2] In view of Remark 3.3, our nomenclature matches that in quantum mechanics. For example, the PDEs (15) are said to be exactly solvable if the corresponding Green’s functions or “propagators” [54, Ch. 2.6] can be determined in closed form. [8], i.e., without integrals or series sum. This is the case for classical SB where $\kappa\equiv\kappa_{0}$, the heat kernel in (4).
@@
$\bm{M}_{t_{1}t_{0}}:=\!\int_{t_{0}}^{t_{1}}\!\!\bm{\Phi}_{t_{1}\tau}\bm{B}(\tau)\bm{B}^{\top}(\tau)\bm{\Phi}_{t_{1}\tau}^{\top}{\rm{d}}\tau\succ\bm{0}.$
In this case, the Markov kernel $\kappa\equiv\kappa_{\rm{Linear}}$ is [61, Sec. III]
$\kappa_{\rm{Linear}}\left(t_{0},\bm{x}_{0},t_{1},\bm{x}_{1}\right)=$ $\left(4\pi(t_{1}-t_{0})\right)^{-n/2}\det\left(\bm{M}_{t_{1}t_{0}}\right)^{-1/2}$
(18) $\exp\left(-\dfrac{\left(\bm{\Phi}_{t_{1}t_{0}}\bm{x}_{0}-\bm{x}_{1}\right)^{\top}\bm{M}_{t_{1}t_{0}}^{-1}\left(\bm{\Phi}_{t_{1}t_{0}}\bm{x}_{0}-\bm{x}_{1}\right)}{4(t_{1}-t_{0})}\right).$
A closed form $\kappa$ such as (4) or (18) helps implement the forward and backward pass in dynamic Sinkhorn recursion (Fig. 2) via simple matrix-vector multiplications. This obviates the need for Feynman-Kac path integrals or other nontrivial solvers [12, 62] which incur function approximation errors.
@@ 4.1 Kernel for $\bm{Q}\succ\bm{0}$
(27) $\kappa_{++}(t_{0},\bm{y},t,\bm{z})=\frac{(\det(\bm{M}))^{1/4}}{(2\pi)^{n/2}\sqrt{\det(\sinh(2\left(t-t_{0}\right)\sqrt{\bm{D}}))}}\exp{\left(-\frac{1}{2}\begin{pmatrix}\bm{y}^{\top}&\bm{z}^{\top}\end{pmatrix}\bm{M}\begin{pmatrix}\bm{y}\\ \bm{z}\end{pmatrix}\right)},$
@@
(28) $\bm{M}=\begin{bmatrix}\sqrt{\bm{D}}\coth{\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)}&-\sqrt{\bm{D}}\>{\rm{csch}}\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)\\ -\sqrt{\bm{D}}\>{\rm{csch}}\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)&\sqrt{\bm{D}}\coth{\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)}\end{bmatrix}.$
@@
The next Proposition (proof in Appendix A.6) states that the matrix $\bm{M}$ in (28) can be seen as positive diagonal scaling of a symplectic positive definite matrix, and is therefore positive definite. This will be useful in establishing properties of $\widehat{\varphi}(t,\bm{x})$ in Proposition 5.1.
Proposition 4.5 (Properties of matrix $\bm{M}$).
For $0\leq t_{0}<t<\infty$, and given positive diagonal matrix $\bm{D}$, the matrix $\bm{M}$ in (28) satisfies
(i) decomposition: $\bm{M}=\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}\bm{M}_{1}\bm{M}_{2}\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}$ where
${\small{\bm{M}_{1}\!=\!\!\begin{bmatrix}\!\cosh(2(t-t_{0})\sqrt{\bm{D}})&-\bm{I}\\ -\bm{I}&\cosh(2(t-t_{0})\sqrt{\bm{D}})\!\end{bmatrix}\!,\bm{M}_{2}\!=\!\!\begin{bmatrix}\!{\rm{csch}}(2(t-t_{0})\sqrt{\bm{D}})&\bm{0}\\ \bm{0}&{\rm{csch}}(2(t-t_{0})\sqrt{\bm{D}})\!\end{bmatrix}\!}},$
(ii) $\bm{M}_{1}\bm{M}_{2}\in{\rm{Sp}}_{++}\left(2n,\mathbb{R}\right)$,
(iii) $\bm{M}\in{\rm{Sym}}_{++}\left(2n,\mathbb{R}\right)$.

### v4

@@ Abstract
Schrödinger bridge is a diffusion process that steers a given distribution to another in a prescribed time while minimizing the effort to do so. It can be seen as the stochastic dynamical version of the optimal mass transport, and has growing applications in generative diffusion models and stochastic optimal control. We say a Schrödinger bridge is “exactly solvable” if the associated uncontrolled Markov kernel is available in closed form, since then the bridge can be numerically computed using dynamic Sinkhorn recursion for arbitrary endpoint distributions with finite second moments. In this work, we propose a regularized variant of the Schrödinger bridge with a quadratic state cost-to-go that incentivizes the optimal sample paths to stay close to a nominal level. Unlike the conventional Schrödinger bridge, the regularization induces a state-dependent rate of killing and creation of probability mass, and its solution requires determining the Markov kernel of a reaction-diffusion partial differential equation. We derive this Markov kernel in closed form, showing that the regularized Schrödinger bridge is exactly solvable, even for non-Gaussian endpoints. This advances the state-of-the-art because closed form Markov kernel for the regularized Schrödinger bridge is available in existing literature only for Gaussian endpoints. Our solution recovers the heat kernel in the vanishing regularization (i.e., diffusion without reaction) limit, thereby recovering the solution of the conventional Schrödinger bridge as a special case. We deduce properties of the new kernel and explain its connections with certain exactly solvable models in quantum mechanics.
@@
Schrödinger bridge, Markov kernel, reaction-diffusion PDE, stochastic optimal control, Hermite polynomial.
@@ 1 Introduction
This solution structure was discovered in [59, 60] and later found to hold in more general settings too, e.g., when prior physics is of the form [^1] Here $\bm{f},\bm{g}$ are regular enough so that strong solutions exist for ${\rm{d}}\bm{x}_{t}=\bm{f}(t,\bm{x}_{t}){\rm{d}}t+\sqrt{2}\bm{g}(t,\bm{x}_{t}){\rm{d}}\bm{w}_{t}$.: ${\rm{d}}\bm{x}_{t}=\bm{f}(t,\bm{x}_{t}){\rm{d}}t+\sqrt{2}\bm{g}(t,\bm{x}_{t}){\rm{d}}\bm{w}_{t}$ generalizing Schrödinger’s prior model $\bm{f}=\bm{0},\bm{g}=$ identity. In such generality, one needs to replace the heat kernel $\kappa_{0}$ with the corresponding $(\bm{f},\bm{g})$-dependent Markov kernels associated with the forward and backward Kolmogorov PDEs, which are advection-diffusion PDEs.
@@
Classical SB has been generalized in multiple directions, e.g., with more general drifts and diffusions [45, 12, 50, 46, 64, 52, 51, 67] and with additional bounded domain constraints [11, 27]. However, works on SB with state cost-to-go are limited [68, 24], [49, Ch. 2.6, 2.7]. In particular, investigating SB with quadratic state cost-to-go and deducing closed form kernels for the same, as done here, are novel.
We emphasize here that by finding the Markov kernel $\kappa$ in closed form, our results enable the numerical solution of SB with quadratic state cost for generic endpoint distributions $\mu_{0},\mu_{1}$ with finite second moments. The solution of SB with quadratic state cost in the special case of Gaussian $\mu_{0},\mu_{1}$ has appeared before in [14, 18]. Different from these prior works, the closed form Markov kernel derived here makes no parametric assumptions on $\mu_{0},\mu_{1}$.
@@ 2 Notations and preliminaries
We will use the following result, which is sometimes referred to as the “central identity” of quantum field theory (QFT) [70, Appendix A].
@@
[72, p. 2][70, p. 15] For a suitably smooth function $f:\mathbb{R}^{n}\mapsto\mathbb{R}$, and matrix $\bm{A}\in{\rm{Sym}}_{++}(n)$,
(9) $\displaystyle\int_{\mathbb{R}^{n}}\!\!\exp\!\left(\!-\frac{1}{2}\bm{x}^{\top}\bm{A}\bm{x}\!\right)\!f(\bm{x}){\rm{d}}\bm{x}=\sqrt{\frac{(2\pi)^{n}}{\det(\bm{A})}}\exp\left(\frac{1}{2}\nabla_{\bm{x}}^{\top}\bm{A}^{-1}\nabla_{\bm{x}}\right)f(\bm{x})\bigg{|}_{\bm{x}=\bm{0}},$
where the exponential of a differential operator is understood as a power series. In particular, for $f(\bm{x})=\exp\langle\bm{b},\bm{x}\rangle$, $\bm{b}\in\mathbb{R}^{n}$, we have $\int_{\mathbb{R}^{n}}\!\exp\!\left(\!-\frac{1}{2}\bm{x}^{\top}\bm{A}\bm{x}+\langle\bm{b},\bm{x}\rangle\!\right)\!{\rm{d}}\bm{x}=\sqrt{\frac{(2\pi)^{n}}{\det(\bm{A})}}\exp\left(\frac{1}{2}\bm{b}^{\top}\bm{A}^{-1}\bm{b}\right)$.
@@ 3.1 Schrödinger factors and reaction-diffusion PDEs
The unique solution to (11) can be determined in terms of the harmonic and coharmonic factors $\varphi,\widehat{\varphi}$, a.k.a. Schrödinger factors, as follows (proof in Appendix A.2). Variants of this result have appeared before; see e.g., [14], [19, Sec. 5].
@@
(15) $-{\mathrm{i}}\displaystyle\frac{\partial\widehat{\Psi}}{\partial\tau}=\left(\Delta-q\right)\widehat{\Psi},\qquad{\mathrm{i}}\displaystyle\frac{\partial\Psi}{\partial\tau}=\left(-\Delta+q\right)\Psi,\qquad{\mathrm{i}}:=\sqrt{-1}.$
The PDEs (15) can be obtained from (12) under imaginary time change (Wick’s rotation) $t\mapsto{\mathrm{i}}\tau$, where $q(\cdot)$ plays the role of quantum mechanical potential. Then, $\Psi,\widehat{\Psi}$ are the wave function and its adjoint, and their product $\widehat{\Psi}(t,\cdot)\Psi(t,\cdot)$ returns the probability amplitude as in (14a). In other words, the Schrödinger factors $(\widehat{\varphi},\varphi)$ in Theorem 3.2 are classical analogues of the wave functions, and the relation (14a) is the classical analogue of the celebrated Born’s relation [7]. For further discussions relating (12) and (15), and their solutions, see [2, Ch. 1.4, 2.1] .
@@ 3.3 Exactly solvable SBs and closed form kernels
We say that an SB problem is exactly solvable if the corresponding (uncontrolled) Markov kernel $\kappa$ is known in closed form [^2] In view of Remark 3.3, our nomenclature matches that in quantum mechanics. For example, the PDEs (15) are said to be exactly solvable if the corresponding Green’s functions or “propagators” [57, Ch. 2.6] can be determined in closed form. [8], i.e., without integrals or series sum. This is the case for classical SB where $\kappa\equiv\kappa_{0}$, the heat kernel in (4).
@@
$\bm{\Gamma}_{t_{1}t_{0}}:=\!\int_{t_{0}}^{t_{1}}\!\!\bm{\Phi}_{t_{1}\tau}\bm{B}(\tau)\bm{B}^{\top}(\tau)\bm{\Phi}_{t_{1}\tau}^{\top}{\rm{d}}\tau\succ\bm{0}.$
In this case, the Markov kernel $\kappa\equiv\kappa_{\rm{Linear}}$ is [64, Sec. III]
$\kappa_{\rm{Linear}}\left(t_{0},\bm{x}_{0},t_{1},\bm{x}_{1}\right)=$ $\left(4\pi(t_{1}-t_{0})\right)^{-n/2}\det\left(\bm{\Gamma}_{t_{1}t_{0}}\right)^{-1/2}$
(18) $\exp\left(-\dfrac{\left(\bm{\Phi}_{t_{1}t_{0}}\bm{x}_{0}-\bm{x}_{1}\right)^{\top}\bm{\Gamma}_{t_{1}t_{0}}^{-1}\left(\bm{\Phi}_{t_{1}t_{0}}\bm{x}_{0}-\bm{x}_{1}\right)}{4(t_{1}-t_{0})}\right).$
A closed form $\kappa$ such as (4) or (18) helps numerically implement the forward and backward pass in dynamic Sinkhorn recursion (Fig. 2) via simple matrix-vector multiplications. This obviates randomized computation with Feynman-Kac path integrals or other nontrivial solvers [12, 65] which incur function approximation errors.
@@ 4.1 Kernel for $\bm{Q}\succ\bm{0}$
(27) $\kappa_{++}(t_{0},\bm{y},t,\bm{z})=\frac{(\det({\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}))^{1/4}}{(2\pi)^{n/2}\sqrt{\det(\sinh(2\left(t-t_{0}\right)\sqrt{\bm{D}}))}}\exp{\!\left(\!-\frac{1}{2}\begin{pmatrix}\bm{y}^{\top}&\bm{z}^{\top}\end{pmatrix}{\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}\begin{pmatrix}\bm{y}\\ \bm{z}\end{pmatrix}\!\right)},$
@@
(28) ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}=\begin{bmatrix}\sqrt{\bm{D}}\coth{\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)}&-\sqrt{\bm{D}}\>{\rm{csch}}\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)\\ -\sqrt{\bm{D}}\>{\rm{csch}}\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)&\sqrt{\bm{D}}\coth{\left(2\left(t-t_{0}\right)\sqrt{\bm{D}}\right)}\end{bmatrix}.$
@@
The next Proposition (proof in Appendix A.6) states that the matrix ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}$ in (28) can be seen as positive diagonal scaling of a symplectic positive definite matrix, and is therefore positive definite. This will be useful in establishing properties of $\widehat{\varphi}(t,\bm{x})$ in Proposition 5.1.
Proposition 4.5 (Properties of matrix ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}$).
For $0\leq t_{0}<t<\infty$, and given positive diagonal matrix $\bm{D}$, the matrix ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}$ in (28) satisfies
(i) decomposition: ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}=\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}{\color[rgb]{0,0,0}{\bm{M}^{(1)}_{tt_{0}}\bm{M}^{(2)}_{tt_{0}}}}\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}$ where
${\color[rgb]{0,0,0}{\bm{M}^{(1)}_{tt_{0}}}}$ $=\begin{bmatrix}\cosh(2(t-t_{0})\sqrt{\bm{D}})&-\bm{I}\\ -\bm{I}&\cosh(2(t-t_{0})\sqrt{\bm{D}})\end{bmatrix},$
${\color[rgb]{0,0,0}{\bm{M}^{(2)}_{tt_{0}}}}$ $=\begin{bmatrix}{\rm{csch}}(2(t-t_{0})\sqrt{\bm{D}})&\bm{0}\\ \bm{0}&{\rm{csch}}(2(t-t_{0})\sqrt{\bm{D}})\end{bmatrix},$
(ii) ${\color[rgb]{0,0,0}{\bm{M}^{(1)}_{tt_{0}}\bm{M}^{(2)}_{tt_{0}}}}\in{\rm{Sp}}_{++}\left(2n,\mathbb{R}\right)$,
(iii) ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}\in{\rm{Sym}}_{++}\left(2n,\mathbb{R}\right)$.
Since $\bm{M}_{tt_{0}}\succ\bm{0}$, the exponential term in (27) is of the form $\exp\left(-\frac{1}{2}{\rm{dist}}_{tt_{0}}^{2}(\bm{y},\bm{z})\right)$ where ${\rm{dist}}_{tt_{0}}(\bm{y},\bm{z})$ is a distance metric. Our next result (proof in Appendix A.7) provides geometric insight that ${\rm{dist}}^{2}_{tt_{0}}(\bm{y},\bm{z})$ is, in fact, the minimal value of the action integral induced by the Lagrangian $L:\mathcal{T}\mathbb{R}^{n}\mapsto\mathbb{R}$ given by [^3] Recall that the diagonal matrix associated with the eigen-decomposition of $\bm{Q}$ is $2\bm{D}$.
(29) $L(\bm{q},\bm{v}):=\dfrac{1}{2}|\bm{v}|^{2}+\bm{q}^{\top}\left(2\bm{D}\right)\bm{q},\quad\left(\bm{q},\bm{v}\right)\in\mathcal{T}\mathbb{R}^{n}.$
Proposition 4.6 (Induced distance in $\kappa_{++}$).
Given $\bm{y},\bm{z}\in\mathbb{R}^{n}$, the minimum value of the action integral associated with the Lagrangian (29):
(30) ${\rm{dist}}_{tt_{0}}^{2}(\bm{y},\bm{z}):=\underset{\gamma(\cdot)\in\big{\{}\mathcal{C}^{2}\left([t_{0},t]\right)\mid\gamma\left(t_{0}\right)=\bm{y},\gamma\left(t\right)=\bm{z}\big{\}}}{\inf}\displaystyle\int_{t_{0}}^{t}L\left(\gamma(\tau),\dot{\gamma}(\tau)\right){\rm{d}}\tau$
equals to $\begin{pmatrix}\bm{y}^{\top}&\bm{z}^{\top}\end{pmatrix}\bm{M}_{tt_{0}}\begin{pmatrix}\bm{y}\\ \bm{z}\end{pmatrix}$, where $\bm{M}_{tt_{0}}$ is given by (28).
Remark 4.7.
Unlike the Markov kernels (4) and (18), the kernel $\kappa_{++}$ in (27) is not a transition probability density because of the reaction term contributing to state-dependent creation or killing of mass. This can be seen by applying (9) to evaluate $\int_{\mathbb{R}^{2n}}\kappa_{++}(t_{0},\bm{y},t,\bm{z}){\rm{d}}\bm{y}{\rm{d}}\bm{z}=\det\left(\sqrt{\bm{D}}\sinh\left(2(t-t_{0})\sqrt{\bm{D}}\right)\right)^{-1/2}\neq 1$.